OpenStudy (anonymous):

~HELP !! FAST !! I NEED HELP ON THIS !! ♥ Combine like terms. 8 + 6x – 11x + 12 A. 14x + 1x B. 20 + 5x C. 20 – 17x D. 20 – 5x

OpenStudy (anonymous):

The answer is D.

OpenStudy (anonymous):

I suggest you try using the app "Mathway" it is a website as well. It will answer these questions quickly for you.

OpenStudy (rina.r):

8+6x-11x+12 8-12+6x-11x 4 -5x

OpenStudy (pixiedust1):

I love Mathway! You can really use any algebra calculator to solve this. (it will start off by reordering and combining!)

OpenStudy (rina.r):

Or 6x-11x+12-8 -5x +4 so u can choose which way u wnat to take.

OpenStudy (pixiedust1):

@Rina.r - am I missing something? I do not see 4 as one of the options.

OpenStudy (rina.r):

what do u mean taht u dont understand where 4 came from

OpenStudy (rina.r):

*that

OpenStudy (pixiedust1):

Looking at her answer options, I do not see one that is -5x+4

OpenStudy (rina.r):

may be the option could be wrong but that's how u do it try it ur self and will will see b coz when u collect like term togather that's what u get u take either 8 to the other side which will be -8 or bring 12 to the left which will be -12

OpenStudy (anonymous):

Sweetie, its D, I already worked it out and got the answer.

OpenStudy (pixiedust1):

Cool, but it's actually solved like this. (8+6x)−11x)+12 =8+6x+(−11x)+12 =(−5x)+20 =−5x+20

OpenStudy (pixiedust1):

So the answer is actually 20-5x

OpenStudy (pixiedust1):

I did too, and checked my work with a few math calculators.

OpenStudy (anonymous):

So did I! virtual high five! lol

OpenStudy (pixiedust1):

Lol! Totally! :D Have a great night!

OpenStudy (anonymous):

Yes , You Guys R All Correct :)